The Award Plus Company manufactures windows. Its manufacturing plant has the capacity to produce 12,000 windows each month. Current production and sales are 10,000 windows per month. The company normally charges $300 per window. Cost information for the current activity level is as​ follows:

Variable costs that vary with number of units produced

Direct materials

$600,000

Direct manufacturing labour

450,000

Variable costs (for setups, materials handling, quality control, and so on)

that vary with number of batches, 200 batches x $1,500 per batch

300,000

Fixed manufacturing costs

200,000

Fixed marketing costs

50,000

Total costs 1,600,000

Award Plus has just received a special​ one-time-only order for 2,000 windows at $250 per window. Accepting the special order would not affect the​ company's regular business or its fixed costs. Award Plus makes windows for its existing customers in batch sizes of 50 windows ​(200 batches x 50 windows per batch =10,000 ​windows). The special order requires Award Plus to make the windows in 40 batches of 50 windows.

Requirement 1. Should Award Plus accept this special​ order? Show your calculations.

Begin by completing an analysis and start by showing the computation of the​ company's operating income without the special order. Next calculate operating income with the special​ order, and then calculate the differences between the two columns. ​(Complete all answer boxes. For amounts with no​ change, make sure to enter​ "0" in the appropriate cells of the Difference​ column.)

Without

One-Time Only

Special Order

10,000 Windows

Revenues

?

Variable costs:

?

Direct materials

?

Direct manufacturing labour

?

Batch manufacturing costs

?

Fixed costs:

?

Fixed manufacturing costs

?

Fixed marketing costs

?

Total costs

?

Operating income

?
Answer #1
Particulars Without Accepting Special Order After Accepting Special Order   Increase / (Decrease)
Sales Revenue 30,00,000 35,00,000 5,00,000
Variable Costs
Direct Materials 6,00,000 7,20,000 1,20,000
Direct Labour 4,50,000 5,40,000 90,000
Variable Manufacturing Overhead 3,00,000 3,60,000 60,000
Total Variable Costs 13,50,000 16,20,000 2,70,000
Contribution Margin 16,50,000 18,80,000 2,30,000
Fixed Costs
Fixed Manufacturing Costs 2,00,000 2,00,000 0
Marketing Costs 50,000 50,000 0
Total Fixed Costs 2,50,000 2,50,000 0
Operating Income 14,00,000 16,30,000 230,000

Increase in Operating Income = Operating Income afetr Special Order - Operating Income without Spending Order

Increase in Operating Income = 16,30,000 - 14,00,000

Increase in Operating Income = $ 2,30,000

Yes the Company should accept the Special Order as the operating income is increased by $ 230,000.

Notes and Supporting Work

Contribution Margin = Sales - Total Variable Costs

Operating Income = Contribution Margin - Total Fixed Costs

Calculation for Accepting the Sales Order

Sales Revenue = 10,000 Units * $ 300 = $ 30,00,000

Sales Revenue for additional sales of 2,000 Units = 2,000 * $ 250 per Unit = $ 500,000

Total Sales Revenue = $ 35,00,000

Total Variable Costs

Variable Manufacturing Overhead = 240 Batches * $ 1,500 per Batch = $ 360,000

Direct Labour = $ 45 per Unit * 12,000 Units = $ 540,000

Direct Materials = $ 60 per Unit * 12,000 Units = $ 720,000

Additional Information

Direct Materials Cost per Unit = 600,000 / 10,000 = $ 60 per Unit

Direct Labour Cost per Unit = 450,000 / 10,000 = $ 45 per Unit

Calculation for 10,000 Units without special order has already been done in the question.

Increase / (Decrease) = Values after Acceptance of Special Order - Values without Accepting Special Order